OpenStudy (anonymous):

hw to solve integral of sin3xcos5x???

OpenStudy (wasiqss):

integration by parts

OpenStudy (anonymous):

by parts??? by taking whiich one???

OpenStudy (wasiqss):

is it sincubex or sin3x

OpenStudy (anonymous):

by parts??? by taking whiich one???

OpenStudy (anonymous):

let u???

OpenStudy (wasiqss):

take any answer wud b same

OpenStudy (wasiqss):

but better take u to b sin3x

OpenStudy (anonymous):

sin3x(-5sin5x)+cos5x(3cos3x)

OpenStudy (anonymous):

use sin(a)cos(b)=(1/2)*[sin(a+b)+sin(a-b)] =>sin(3x)cos(5x)=(1/2)*[sin(8x)-sin(2x)] which is easy to integrate

OpenStudy (anonymous):

u just need to expand above

OpenStudy (anonymous):

ghass1978, by using half angle???

OpenStudy (anonymous):

sin(a+b)=sin(a)cos(b)+sin(b)cos(a) and sin(a-b)=sin(a)cos(b)-sin(b)cos(a) then sin(a+b)+sin(a-b)=2sin(a)cos(b) so sin(a)cos(b)=(1/2)*[sin(a+b)+sin(a-b)]

OpenStudy (anonymous):

tq 4 helping me.... ^_^